Crawford, Inc., has 105,000 shares of stock outstanding. Each share is worth $72, so the company’s market value of equity is $7,560,000. Suppose the firm issues 20,000 new shares at the following prices: $72, $69, and $64. What will the effect be of each of these alternative offering prices on the existing price per share? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16. Leave no cells blank; if there is no effect select "No change" from the dropdown and enter "0".)
Answer #1

New price = (current price*old shares+new share price*add. shares)/(old shares+add. shares)

=(72*105000+72*20000)/(105000+20000)

=72

New price = (current price*old shares+new share price*add. shares)/(old shares+add. shares)

=(72*105000+69*20000)/(105000+20000)

=71.52

New price = (current price*old shares+new share price*add. shares)/(old shares+add. shares)

=(72*105000+64*20000)/(105000+20000)

=70.72